Circular No. 8445

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004gj IN IC 701 Further to IAUC 8443, J. Burket and W. Li report the LOSS discovery

of an apparent supernova on unfiltered KAIT images taken on Nov. 23.52 (mag ∼ 18.3) and 24.55 UT (mag ∼ 18.4). SN 2004gj is located at α = 11^{h}30^{m}59^{s}.63, δ = +20^{o}28′06′′.8 (equinox 2000.0), which is 14′′.1 west and 2′′.1 south of the nucleus of the brighter eastern component of the interacting galaxy IC 701. A KAIT image taken on May 21.19 (limiting mag ∼ 19.5) showed nothing at this position.

V574 PUPPIS = NOVA PUPPIS 2004 N. N. Samus and E. Kazarovets, Institute of Astronomy, Moscow, in-

form us that the designation V574 Pup has been given to this nova (cf. IAUC 8443).

CCD magnitudes (uncertainty ± 0.05 mag) by G. Sostero, Udine, Italy, on Nov. 22.093 UT: B = 8.48, V = 7.62, R_{c} = 7.06, I_{c} = 6.45. CCD photometry reported by H. Yamaoka, Kyushu University: Nov. 22.716, V = 6.93, B = 7.79 (S. Kiyota, Tsukuba, Ibaraki, Japan); 22.717, R_{c} = 6.25, I_{c} = 5.54 (Kiyota); 22.758, V = 6.98 (K. Nakajima, Mie, Japan); 22.787, I_{c} = 5.85, R_{c} = 6.53 (H. Maehara, Saitama, Japan); 22.788, V = 7.14, B = 7.79 (Maehara); 23.806, I_{c} = 7.18 (Maehara); 23.807, R_{c} = 8.02, V = 8.90 (Maehara); 23.808, B = 9.53 (Maehara).

Visual magnitude estimates (partly via Yamaoka): Nov. 22.068 UT, 7.9 (R. Y. Shida, São Paulo, Brazil); 22.694, 7.1 (K. Hirosawa, Aichi, Japan); 22.825, 7.0 (S. Takahashi, Shiga, Japan); 23.611, 8.6 (A. Pearce, Nedlands, W. Australia); 23.685, 8.8 (Y. Maeda, Nagasaki, Japan); 23.809, 8.9 (Hirosawa); 24.064, 8.8 (P. Schmeer, Bischmisheim, Germany); 24.095, 8.3 (Shida); 24.171, 8.7 (R. J. Bouma, Groningen, The Netherlands); 24.229, 8.9 (A. Kammerer, Malsch, Germany); 24.25, 8.4 (W. Souza, São Paulo, Brazil). Corrigendum. On IAUC 8443, line 4, for Ibaragi-ken, read Ibaraki-ken,

COMETS 161P/HARTLEY-IRAS AND 162P/SIDING SPRING Two additional comets have now been numbered: 161P/Hartley-IRAS

(P/1983 V1 = P/2004 V2; IAUC 8428) and 162P/Siding Spring (P/2004 TU_{12}; IAUC 8436).

2004 November 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green